OpenStudy (anonymous):

If f(x) = 3x – 2 and g(x) = 6 – 4x, find f(x) + g(x).

OpenStudy (caozeyuan):

just sum them up and you are there ! f(x)+g(x)= (3x-4x)+(-2+6)=-x+4

OpenStudy (anonymous):

Just add the two functions !

OpenStudy (anonymous):

@caozeyuan so the answer is???

OpenStudy (caozeyuan):

as I posted: -x+4. BTW, could you teach me how to @ people, as you did there?

OpenStudy (anonymous):

@caozeyuan just press shift and 2 at the same time and then type the users' name

OpenStudy (caozeyuan):

@jnn452 why x=4? it's not a equation, x is a variable!

OpenStudy (anonymous):

sorry, i just saw as they should be equal, so i think you might be right.